OpenStudy (anonymous):

Show that if f is an odd function such that 0 is in the domain of f, then f(0)=0

OpenStudy (zarkon):

\[f(-x)=-f(x)\] \[f(0)=f(-0)=-f(0)\] \[f(0)=-f(0)\Rightarrow f(0)+f(0)=0\] \[2f(0)=0\Rightarrow f(0)=0\]